find three positive consecutive even integers such that the product of the first and second is 8 more than 38 times the third.
Answer by checkley79(3341)

Let x, x+2 & x+4 be the 3 integers.
x(x+2)=38(x+4)+8
x^2+2x=38x+152+8
x^2+2x-38x-160=0
x^2-36x-160=0
(x-40)(x+4)=0
x-40=0
x=40 ans.
x=4=0
x=-4 ans.
Proof:
40*42=38*44+8
1,680=1,672+8
1,680=1,680